the speed of a stream is 3 mph. A boat travels 3 miles upstream in the same time it takes to travel 9 miles downstream. What is the speed of the boat in still water?

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

let speed in still water = x mph
current speed 3 mph
..
upstream speed = x-3 mph
downstream speed = x+3 mph

Distance upstream=3
Distance downstream=9
Time = same
3/(x-3)= 9/(x+3)
3*(x+3)= 9*(x-3)
3x+3=9x-27
9x-3x=27 +3
6x=30
/6
x=5mph

Answer by richwmiller(17219)

r*t=d
(r-3)*t=3
(r-3*t=3
r*t-3*t=3
(r+3)*t=9
r*t+3*t=9
r*t-3*t=3
add
2r*t=12
r*t=12/2
r*t=6
ok
r*t-3*t=3
6-3*t=3
6-3=3*t
3=3*t
3/3=t
1=t
t=1 hour
r*1=6
r=6/1
r=6 mph
check
(r+c)*t=9
(6+3)*1=9
(9)*1=9
9=9
(r-c)*t=3
(6-c)*t=3
(6-3)*1=3
3*1=3
3=3
ok